Quinine (C20H24N2O2) is the most important alkaloid derived from cinchona bark. It is used as an antimalarial drug. For quinine pKb1 = 5.1 and pKb2 = 9.7 (pKb = −log(Kb)). Only 1.0 g of quinine will dissolve in 1900.0 mL of water. Calculate the pH of saturated aqueous solution of quinine. Consider only the reaction Q + H2O equilibrium reaction arrow QH + + OH − described by pKb1, where Q = quinine.
Answer #1

page:oi Answeni. information, According to the given to calculate pH. Giveni preachion Ps, Q+ H2O = QH+ OH since, 21 fs mon opageiol mat: 1962xto? x49u3 x16 = 162% 7,903 x1ő 2 1.2876X108 Da 11 13x164M :: GH] = 3 = 123* 174M thon, pot = -log[on] 5-104